(1/16)^3-y =4^-3(2y-7)
(3-y) is all raised and -3(2y-7) is all raised

Answer by stanbon(75887)

(1/16)^(3-y) =4^(-3(2y-7))
--------------------------
(2^-4)^(3-y) = (2^2)^(-6y+21)
-------
2^(-12+4y) = 2^(-12y+42)
------
-12+4y = -12y + 42
-----
16y = 54
y = 27/8
==================
Cheers,
Stan H.